OpenStudy (anonymous):

Can someone please please please help me! Find the angle between the given vectors to the nearest tenth of a degree. u = <8, 4>, v = <9, -9> 81.6° 25.8° 35.8° 71.6° I keep getting the same answer and it isnt even one of my choices.

OpenStudy (anonymous):

Im supposed to use \[\cos \theta=\frac{ u*v }{ \left| u \left| v \right| \right| }\], right?